OpenStudy (anonymous):

II WILL FAN AND MEDAL!! Solve for f. seven-twelfths times f equals three-fifths

OpenStudy (help_people):

Please stop saying you will fan and medal if you will not. You did not fan or medal for the last question so....

OpenStudy (anonymous):

Dividing by a fraction is multiplying by the reciprocal. f=35⋅127

OpenStudy (anonymous):

yes i did i did it for someone else

OpenStudy (help_people):

is there any multiple chpoice answers?

OpenStudy (anonymous):

answer is both 1 and 1/35

OpenStudy (anonymous):

yes

Nnesha (nnesha):

\[\huge\rm \frac{ 7 }{ 12 }f = \frac{ 3 }{ 5 }\] again cross multiply

OpenStudy (anonymous):

Solve for f. seven-twelfths times f equals three-fifths Choose one answer. A. Thirty-five thirty-sixths B. One and thirty-five thirty-sixths C. Twenty-one sixtieths D. One and one thirty-fifth

OpenStudy (anonymous):

it is a

OpenStudy (help_people):

ok thank you that is better, lets start off saying which one do you think it is?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

byt he said A

OpenStudy (anonymous):

so i guess it is A

OpenStudy (help_people):

Well if you think it is b, then can you show work for why you think it is b?

OpenStudy (help_people):

@095ut89h you still here?

OpenStudy (help_people):

is this the equation? \[7/12f \times3/5\]

OpenStudy (anonymous):

it could be A or B i don't know

OpenStudy (help_people):

Well lets slow down can you show me how it ovule be A or B, if you are guessing then....

OpenStudy (help_people):

*could

Nnesha (nnesha):

\[\huge\rm \frac{ 7 }{ 12 }f = \frac{ 3 }{ 5 }\] again cross multiply  cross multiply like in this example